# On the variety of general position problems under vertex and edge removal

### v1

@@ 4 Edge removal
[…]
@@
Let $X=\{w_{1},\dots,w_{n}\}$. It is straightforward to check that $X$ is a general position set of $H_{n}$. Since there is a unique shortest path between $x$ and $y$ that does not contain any vertex from $X$, it follows from Theorem 2.2 (iii) that GXG-X is convex, thus proving that $X$ is the largest dual general position set of $H_{n}$. Hence ${\rm gp}_{\rm d}(H_{n})\geq n$. On the other hand, removing the edge $e=xy$, every edge of $H_{n}-e$ is the middle edge of some isometric path $P_{4}$. In view of [17, Proposition 3.3] we have ${\rm gp}_{\rm d}(H_{n}-e)=0$. $\square$
@@ Acknowledgements
This work was supported by the Slovenian Research and Innovation Agency (ARIS) under the grants P1-0297, N1-0355, and N1-0285.

### v2

@@ 4 Edge removal
For the sharpness, note first that $s(K_{n})=n$ and $s(K_{n}-e)=2$, hence the lower bound is sharp. For the upper bound, let $X_{n}$, $n\geq 2$, be the graph obtained from the disjoint union of two copies of $K_{n}$, say $K$ and $K^{\prime}$ and a vertex $x$, by adding an edge $e=uv$ between $K$ and $K^{\prime}$, and joining $x$ to a vertex $u^{\prime}$ in $K$ and a vertex $v^{\prime}$ in $K^{\prime}$, where $\{u,v\}\cap\{u^{\prime},v^{\prime}\}=\emptyset$. Then we infer that ${\rm gp}_{\rm t}(X_{n})=2(n-2)$ and ${\rm gp}_{\rm t}(X_{n}-e)=2(n-1)$. $\square$
@@
The sharpness of the upper bound in Theorem 4.2 is demonstrated by graphs that contain bridges. Consider next the graphs $Z_{n}$, $n\geq 3$, constructed as follows. Take two disjoint complete graphs $K_{n}$, add an edge $e$ between them, and add a path of length $3$ between them disjoint from $e$. Then we can infer that ${\rm gp}_{\rm o}(Z_{n})=2n-4$ and that ${\rm gp}_{\rm o}(Z_{n}-e)=2n-2$. In the case $n=3$, we thus also have a graph without bridges for which the upper bound in Theorem 4.2 is sharp. It remains to be seen whether the upper bound in Theorem 4.2 is sharp also on graphs without bridges with an arbitrary large other general position number.
@@
Let $X=\{w_{1},\dots,w_{n}\}$. It is straightforward to check that $X$ is a general position set of $H_{n}$. Since there is a unique shortest path between $x$ and $y$ that does not contain any vertex from $X$, it follows that GXG-X is convex. In addition, since $\{x,y,v_{1},v_{2},v_{3},v_{4}\}$ induces a convex $C_{6}$, using Theorem 2.2(iii) we infer that none of these vertices lies in a dual general position set. We can conclude that $X$ is the largest dual general position set of $H_{n}$, therefore ${\rm gp}_{\rm d}(H_{n})=n$. On the other hand, removing the edge $e=xy$, every edge of $H_{n}-e$ is the middle edge of some isometric path $P_{4}$. In view of [18, Proposition 3.3] we have ${\rm gp}_{\rm d}(H_{n}-e)=0$. $\square$
@@ Acknowledgements
This work was supported by the Slovenian Research and Innovation Agency (ARIS) under the grants P1-0297, N1-0355, N1-0285, N1-0431, and J1-70045.